### v3

# Sommerfeld Enhancement for Vector Dark Matter in the Fundamental Representation of $SU(2)_{L}$

Abstract
In this work, we study an extension of the Standard Model which include a new massive vector field in the fundamental representation of $SU(2)_{L}$. The neutral component of the new vector field is a natural dark matter candidate. We compute the Annihilation Cross Section with Sommerfeld Enhancement and the Photon Flux arising from Dark Matter annihilation events. We found that the model exhibits a resonant mass part around the range of $2-10$ TeV. This resonance is influenced by the choice of portal Higgs coupling parameter. We show that these prediction can be partially tested by CTA in the near future.
1 Introduction
Currently, one of the most urgent problems in Particle Physics revolves around the elucidation of Dark Matter’s (DM) fundamental nature  [1]. It is well known that, to account for the overarching phenomenology of DM, it suffices to posit the existence of a new particle falling within a mass range spanning from a few GeV to a few TeV, characterized by an interaction strength with the standard sector on the order of the weak interaction. This phenomenon is commonly referred to as the ”WIMP miracle”, with WIMP representing ”Weakly Interacting Massive Particle” [2].
An intriguing realization of the WIMP idea is to postulate the DM as the neutral component of an electroweak multiplet. This paradigm, usually called “Minimal Dark Matter”, was proposed long ago [3] and have been studied in detail for scalar field [4] and fermionic fields [5].
In the case of vector DM (VDM), an effective model including a VDM candidate originated from a vector field in the adjoint representation of $SU(2)_{L}$ was first proposed in [6] while an ultraviolet complete model was developed in [7]. On the other hand, models including a vector DM candidate coming from an electroweak doublet were studied in  [8, 9]. These vector models are consistent with experimental data, if the new fields has mass in the range of some few TeV. In this mass range, direct detection [10] becomes difficult and the discovery at colliders seems to be challenging even for future colliders [6].
In this circumstances, indirect search can be a privileged alternative to discover or rule out these models. However, it is well known that to obtain an accurate value of the DM annihilation cross section in the non-relativistic regime, it is necessary to take into account non-perturbative effects like the so called Sommerfeld Enhancement (SE).
The SE effect was computed for the cases of scalar DM [4, 11, 12, 13, 14], fermion DM field [5, 15] in the doublet and triplet representation of $SU(2)_{L}$ and vector DM in the adjoint representation  [7]. In this work, we focus on computing the SE for the annihilation of VDM in the fundamental representation of $SU(2)_{L}$ and the photon flux (relevant for Indirect Detection  [10, 16, 17, 18]) produced by pair annihilation of the VDM. As we will see in the following section, a unique feature of our case is the presence of non-minimal interactions between the massive vector doublet and the standard gauge bosons.
Our work is organized as follows. In section 2, we describe the main aspects of the model. Section 3 is devoted to the computation of the electroweak, annihilation matrices and annihilation cross section with SE. In section 4 we apply our results to compute the photon flux from DM annihilation at the center of the Milky Way and describe our results. Finally, in section 5 we state our conclusions.
2 The Model
We start by recalling the main characteristics of the model initially developed in reference  [8]. In this context, the Standard Model (SM) is extended by the introduction of a vector field $V_{\mu}$ which transforms as (1, 2, 1/2) under the standard group $SU(3)_{c}\times SU(2)_{L}\times U(1)_{Y}$. This new vector field can be explicitly written as:
$V_{\mu}=\begin{pmatrix}V_{\mu}^{+}\\ V_{\mu}^{0}\end{pmatrix}=\begin{pmatrix}V_{\mu}^{+}\\ \dfrac{V_{\mu}^{1}+i~{}V_{\mu}^{2}}{\sqrt{2}}\end{pmatrix}$ (1)
In principle, the two neutral states $V_{\mu}^{1}$ and $V_{\mu}^{2}$ (the vector and axial-vector fields) can have different masses. However, unitarity constrains force the neutral components $V_{\mu}^{1}$ and $V_{\mu}^{2}$ to be fully degenerate  [8]. In this case they combine into a complex neutral field $V_{\mu}^{0}$.
The Lagrangian of the model can be written as:
$\mathcal{L}=\mathcal{L}_{SM}$ $-\dfrac{1}{2}\left(D_{\mu}V_{\nu}-D_{\nu}V_{\mu}\right)^{\dagger}\left(D^{\mu}V^{\nu}-D^{\nu}V^{\mu}\right)+m_{V}^{2}V_{\nu}^{\dagger}V^{\nu}$
$-\lambda_{2}(\phi^{\dagger}\phi)(V_{\mu}^{\dagger}V^{\mu})-\lambda_{3}(\phi^{\dagger}V_{\mu})(V^{\mu\dagger}\phi)$
$+\lambda_{4}\left[(\phi^{\dagger}V_{\mu})(\phi^{\dagger}V^{\mu})+(V^{\mu\dagger}\phi)(V_{\mu}^{\dagger}\phi)\right]$
$-\alpha_{2}(V_{\mu}^{\dagger}V^{\mu})(V_{\nu}^{\dagger}V^{\nu})-\alpha_{3}(V_{\mu}^{\dagger}V^{\nu})(V_{\nu}^{\dagger}V^{\mu})$
$+i\dfrac{g^{\prime}}{2}\kappa_{1}V_{\mu}^{\dagger}B^{\mu\nu}V_{\nu}+ig\kappa_{2}V_{\mu}^{\dagger}W^{\mu\nu}V_{\nu}$ (2)
where $B^{\mu\nu}$ represents the Abelian strength field of $U(1)_{Y}$ and $W^{\mu\nu}=W^{\mu\nu a}\frac{\tau^{a}}{2}$ is the non-Abelian strength field of $SU(2)_{L}$. The lagrangian mass of the vector field $\tilde{m}$ will be corrected after the Higgs boson acquires its vacuum expectation value, due to the interaction terms between the Higgs field and the vector field present in the potential. We will call $m$ the physical mass of the vector boson.
Notice that the last two terms of Lagrangian (2) describe non-minimal interactions between the new vector field and the gauge bosons. The appearance of this non-minimal dimension 4 terms is a special feature of having introduced a vector doublet. Similar non-minimal interactions in the scalar and fermion cases happen to be higher dimensional. In order to avoid a direct photon interaction with $V_{\mu}^{0}$, the otherwise arbitrary constants $\kappa_{1}$ and $\kappa_{2}$ have to be equal. For the purpose of comparing, we work with two values of $\kappa=\kappa_{1}=\kappa_{2}$: $\kappa=0$ and $\kappa=1$. The last value coincides with the choice made in [8].
Another important sector of the Lagrangian is the part of the potential which describes the interaction between the Higgs field and our massive vector doublet. Of course, the coupling constant that regulates the interactions between the Higgs field and the new vector field are free parameters. However, there is no reason to think that an important hierarchy arise among these couplings. Consequently, and for the sake of simplicity, we will take the assumption that all of them are of the same order $\lambda_{2}\approx\lambda_{3}\approx\lambda_{4}\approx\lambda$. In this way, $\lambda$ will be the only constant regulating the Higgs portal. After the electroweak symmetry breaking, the physical mass of $V_{\mu}$ ($M_{V}$) receives contributions from the Proca term ($+m_{V}^{2}V_{\nu}^{\dagger}V^{\nu}$) but also from the interactions with the Higgs field.
Finally, the $V_{\mu}$ quartic self-interactions will not contribute significantly to the SE since this phenomenon is dominated by interactions mediated by particles much lighter than $V_{\mu}$, like the electroweak gauge bosons and the Higgs boson. Consequently, we can set $\alpha_{2}=\alpha_{3}=0$.
Due to a fortunate conspiracy between the Lorentz and gauge structures of the model, the new vector field can only appear in pairs in the Lagrangian. That means that the Lagrangian is invariant by a $Z_{2}$ transformation under which the new vector is odd and all the standard fields are even. Consequently, this accidental $Z_{2}$ symmetry makes the lighter component of $V_{\mu}$ stable. This stable component happens to be the neutral one ($V_{\mu}^{0}$)
3 Sommerfeld Enhancement Computation
3.1 Basic Formalism
The so called “Sommerfeld enhancement” is a non-perturbative effect due to the resumation of initial states interactions. This effect may produce sensible consequences for the annihilation cross section in the non-relativistic regime [5, 19]. In the next few paragraphs we will review the formalism used to compute the Sommerfeld enhancement effect [20] and the we will apply it to our model.
Figure 1: Sommerfeld enhancement effect
3.1.1 The Two-Particle State and Schrödinger Equation
To describe the state of two DM-DM particles, which must be neutral in charge and CP-even, we consider the following combinations:
|$\ket{g}=\begin{pmatrix}v_{0}v_{0}\\ a_{0}a_{0}\\ v_{0}a_{0}\\ v_{+}v_{-}\end{pmatrix}$ (3)
The order of the components in the vector of equation (3) is arbitrary, and we can use different orderings for the pairs of these doublets.
Since this effect is non-relativistic, we must use the Schrödinger equation for two bodies, derived from the Schwinger-Dyson equation [5], which is:
$\left[-\dfrac{1}{M_{V}}\dfrac{d^{2}}{dr^{2}}-\frac{l(l+1)}{M_{V}r^{2}}-E+V(r)-\dfrac{iN\Gamma\delta(r)}{8\pi r^{2}}\right]g_{ab}(r)=0$ (4)
where index $a$ represents the initial state of the process and $b$ the final state and $r$ is the relative position. The wave function $g_{ab}$ describes the two particles in a doublet of $SU(2)_{L}$. As we have seen, the potential includes interactions with gauge and Higgs bosons, corresponding to Coulomb (long-range) and Yukawa (short-range) potentials. The kinetic energy of these particles is $E\approx\frac{M_{V}v^{2}}{4}$, where $v$ is the relative velocity between them. The $\Gamma$ term is the imaginary potential and corresponds to an absorptive part, arising in the low-energy limit of quartic interactions. Finally, $N$ is the number of spin states, given by $N=(2l+1)^{2}$, which in our case is 9.
We can consider polarized initial states in the general case, but the main contribution in the non-relativistic limit occurs in the s-wave, i.e., when $l=0$, where the total angular momentum is given solely by the spin contribution. Therefore, the Schrödinger equation is:
$\left[-\dfrac{1}{M_{V}}\dfrac{d^{2}}{dr^{2}}-E+V(r)-\dfrac{iN\Gamma\delta(r)}{8\pi r^{2}}\right]g_{ab}(r)=0$ (5)
To solve this equation (5), we need the boundary conditions given by [5]:
$g_{ab}(r\rightarrow 0)=\delta_{ab}$ (6)
$g_{ab}(r\rightarrow\infty)=d_{ab}(E)e^{i\sqrt{M_{V}E}r}e^{i\left(\frac{M_{V}\alpha}{2M_{V}E}\ln(2\sqrt{M_{V}E}r)\right)}$ (7)
where $\delta_{ab}$ is a Kronecker delta, $\alpha$ is the fine-structure constant, and $d_{ab}$ are the Sommerfeld enhancement factors. The asymptotic condition of (7) is valid for large distances encompassing the Coulomb and Yukawa potentials. The equation (5) can be solved numerically as it does not have an analytical solution. Once the solutions are determined, we use the optical theorem, which relates the imaginary part of the Green’s functions to the thermally averaged total annihilation cross-section [5, 7] of DM particles:
$\braket{\sigma v}_{XX^{\prime}}=c\sum_{a,b}\sum_{J,J_{z}}(\Gamma_{XX^{\prime}}^{J,J_{z}})_{ab}~{}d_{3a}(E)d_{3b}^{\ast}(E)$ (8)
where $c$ is a normalization constant (2 for $a=b$, 1 for $a\neq b$). The fixed index 3 is related to the definition of the vector $\vec{s}$ (14), to select dark matter annihilation $(V_{0}V_{0}\longrightarrow XX^{\prime})$ with $X$ and $X^{\prime}$ being particles of the Standard Model.
3.1.2 Non-Relativistic Limit
To ensure that the dark matter vector remains non-relativistic, we must carry out the following non-relativistic limit of the operators, as defined in [7].
$V_{\mu}^{0}(x)=\dfrac{1}{\sqrt{2M_{V}}}[v_{0}(\vec{x})\epsilon_{\mu}(x)e^{iM_{V}t}+a_{0}^{\dagger}(\vec{x})\epsilon_{\mu}^{\ast}(x)e^{-iM_{V}t}]\hskip 14.22636pt$ (9)
$(V_{\mu}^{0})^{\ast}(x)=\dfrac{1}{\sqrt{2M_{V}}}[a_{0}(\vec{x})\epsilon_{\mu}(x)e^{iM_{V}t}+v_{0}^{\dagger}(\vec{x})\epsilon_{\mu}^{\ast}(x)e^{-iM_{V}t}]\hskip 14.22636pt$ (10)
$V_{\mu}^{+}(x)=\dfrac{1}{\sqrt{2M_{V}}}[v_{+}(\vec{x})\epsilon_{\mu}(x)e^{iM_{V}t}+v_{-}^{\dagger}(\vec{x})\epsilon_{\mu}^{\ast}(x)e^{-iM_{V}t}]\hskip 14.22636pt$ (11)
$V_{\mu}^{-}(x)=\dfrac{1}{\sqrt{2M_{V}}}[v_{+}^{\dagger}(\vec{x})\epsilon_{\mu}^{\ast}(x)e^{-iM_{V}t}+v_{-}(\vec{x})\epsilon_{\mu}(x)e^{iM_{V}t}]\hskip 14.22636pt$ (12)
Here, $v_{0},a_{0},v_{+},v_{-}$ represent the non-relativistic limits of the dark matter vector field. $m$ denotes the mass of the neutral component of the dark matter vector field, and $\epsilon_{\mu}$ signifies the polarization vector. These polarization vectors adhere to the transverse conditions and an orthonormal relation:
$p^{\mu}\epsilon_{\mu}=0;\hskip 28.45274pt\epsilon_{\mu}\epsilon_{\nu}=-\delta_{\mu\nu}$ (13)
3.2 Electroweak Matrix
3.2.1 Gauge Bosons-Dark Matter Interaction
We aim to investigate the annihilation of dark matter particles into standard model ones. Specifically, we consider the pairs ($v_{0}v_{0}$), for which such annihilation occurs. The most important contributions come from processes involving the exchange of $\gamma$, $Z$, $W^{\pm}$ bosons, having in their initial states ($v_{0}v_{0}$) , ($v_{0}a_{0}$), or ($v_{+}v_{-}$) pairs. Our approach involves deriving an effective action that accurately describes these pairs. For this purpose, we define the following vector:
$s(x,\vec{y})=\begin{pmatrix}\dfrac{v_{0}(x)v_{0}(x^{0},\vec{y})}{\sqrt{2}}\\ \dfrac{a_{0}(x)a_{0}(x^{0},\vec{y})}{\sqrt{2}}\\ v_{0}(x)a_{0}(x^{0},\vec{y})\\ v_{+}(x)v_{-}(x^{0},\vec{y}).\end{pmatrix}$ (14)
The disparity in the normalization of the first component in (14) arises from the fact that they pertain to states composed of two identical particles. The effective action is as follows:
$S_{eff,1}=-\int$ $\dfrac{d^{4}xd^{3}y}{8\pi|\vec{x}-\vec{y}|}(J_{A}^{0}(x)J_{A}^{0}(x^{0},\vec{y})$
$+J_{Z}^{0}(x)J_{Z}^{0}(x^{0},\vec{y})e^{-M_{Z}|\vec{x}-\vec{y}|}$
$+2J_{W_{+}}^{0}(x)J_{W_{-}}^{0}(x^{0},\vec{y})e^{-M_{W}|\vec{x}-\vec{y}|})$ (15)
The currents are calculated from the cubic interactions of the Lagrangian (2) between dark matter and gauge bosons. From these expressions, we employ the non-relativistic limit to obtain the electroweak matrix, which we will use to solve the Schrödinger equation (5).
In the non-relativistic limit, the first term of equation (3.2.1) characterizes the Coulomb potential between non-relativistic particles, whereas the remaining terms depict the Yukawa potentials related to the exchange of $Z$ and $W$ bosons. Upon employing vector (14), the effective action transforms into the following expression:
$S_{eff,1_{NR}}=-\int d^{4}xd^{3}y~{}s(x,\vec{y})^{\dagger}V_{g-V}(|\vec{x}-\vec{y}|)s(x,\vec{y})$ (16)
where $V_{g-V}(|\vec{x}-\vec{y}|)$ is electroweak potential matrix (interaction gauge boson-vector dark matter), with $r=|\vec{x}-\vec{y}|$;
$V_{g-V}(r)=-\dfrac{g^{2}}{8\pi r}\begin{pmatrix}4a^{2}e^{-M_{Z}r}&0&0&0\\ 0&4a^{2}e^{-M_{Z}r}&0&0\\ 0&0&16a^{2}e^{-M_{Z}r}&b^{2}e^{-M_{W}r}\\ 0&0&b^{2}e^{-M_{W}r}&8a^{2}e^{-M_{Z}r}-2(c_{W}(1-(\kappa_{1}+\kappa{2})-s_{W}t_{W}^{2})^{2}\end{pmatrix}$ (17)
with $a=(c_{W}(\kappa_{1}+\kappa_{2})+s_{W})$, $b=(\frac{3}{8}-2\sqrt{2})(\kappa_{1}+\kappa_{2})+2\sqrt{2}$ In appendix A we show how to compute this matrix.
3.2.2 Higgs Boson-Dark Matter Interaction
We need to take into account the interaction between the Higgs Boson and the Dark Matter vector. Our primary focus lies in understanding the cubic interaction, as it is crucial for unraveling the nature of this interaction. For this purpose, we parameterize the Higgs doublet as follows:
$\phi(x)=\begin{pmatrix}\phi^{+}(x)\\ \dfrac{v+h(x)}{\sqrt{2}}+i\phi^{0}(x)\end{pmatrix}$ (18)
where $\phi^{+}$ and $\phi^{0}$ are would-be Goldtone bosons, and $v$ is vacuum expectation value.
Now, when we substitute $V_{\mu}$, $V_{\mu}^{\dagger}$, $\phi$, and $\phi^{\dagger}$ into the Lagrangian (2), we can derive the electroweak matrix describing the interaction between the Higgs and the Dark Matter vector, which takes the form:
$V_{h-V}(r)=\dfrac{v^{2}}{32\pi M_{V}^{2}r}\begin{pmatrix}8\lambda_{4}^{2}a&0&0&0\\ 0&8\lambda_{4}^{2}a&0&0\\ 0&0&\dfrac{1}{2}(\lambda_{2}+\lambda_{3})^{2}a&(\lambda_{3}+2\lambda_{4})^{2}e^{-M_{W}r}\\ 0&0&(\lambda_{3}+2\lambda_{4})^{2}e^{-M_{W}r}&\lambda_{2}^{2}a\end{pmatrix}$ (19)
with $a=(2e^{-M_{h}r}+e^{-M_{Z}r})$. In Appendix A, we illustrate the methodology for computing this matrix.
3.3 Annihilation Matrix
3.3.1 Gauge Boson-Dark Matter Vector
When Dark Matter particles annihilate into Standard Model ones, it becomes essential to compute the annihilation matrix, specifically examining the interaction between Gauge Bosons and the Dark Matter. This requires the use of the quartic terms from Lagrangian (2). The effective action for these quartic terms, in the non-relativistic limit, can be expressed as follows:
$S_{eff,2NR}$ $=2i\int~{}d^{4}xd^{3}y~{}s(x,\vec{y})^{\dagger}\Gamma_{gauge}\delta^{3}(|\vec{x}-\vec{y}|)s(x,\vec{y})$ (20)
In Appendix B, the computation of these terms is detailed. Considering the different channels and total spin states, we obtain the following annihilation matrices:
$\Gamma_{WW}^{J=0}=\dfrac{2g^{4}}{64\sqrt{3}\pi M_{V}^{2}}\begin{pmatrix}2&0&0&0\\ 0&2&0&0\\ 0&0&1&1\\ 0&0&1&1\end{pmatrix}$ (21)
$\Gamma_{WW}^{J=2}=\dfrac{g^{4}}{64\pi M_{V}^{2}}\left(\dfrac{1+3\cos(2\theta)}{2\sqrt{6}}+\sin^{2}\theta\right)\begin{pmatrix}2&0&0&0\\ 0&2&0&0\\ 0&0&1&1\\ 0&0&1&1\end{pmatrix}$ (22)
$\Gamma_{AA}^{J=0}=\dfrac{e^{4}}{128\sqrt{3}\pi M_{V}^{2}}\begin{pmatrix}2a^{2}&0&0&0\\ 0&2a^{2}&0&0\\ 0&0&a^{2}&4a^{2}\\ 0&0&4a^{2}&4\end{pmatrix}$ (23)
$\Gamma_{AA}^{J=2}=\dfrac{e^{4}}{256\pi M_{V}^{2}}\left(\dfrac{1+3\cos(2\theta)}{2\sqrt{6}}+\sin^{2}\theta\right)\begin{pmatrix}2a^{2}&0&0&0\\ 0&2a^{2}&0&0\\ 0&0&a^{2}&4a^{2}\\ 0&0&4a^{2}&4\end{pmatrix}$ (24)
with $a=(1-\tan\theta_{W})$
$\Gamma_{ZZ}^{J=0}$ $=\dfrac{g^{4}}{128\sqrt{3}\pi M_{V}^{2}}\begin{pmatrix}2b^{2}&0&0&0\\ 0&2b^{2}&0&0\\ 0&0&2b^{2}&-4c\\ 0&0&-4c&4\sin^{4}\theta_{W}\end{pmatrix}$ (25)
$\Gamma_{ZZ}^{J=2}$ $=\dfrac{g^{4}}{256\pi M_{V}^{2}}\left(\dfrac{1+3\cos(2\theta)}{2\sqrt{6}}+\sin^{2}\theta\right)\begin{pmatrix}2b^{2}&0&0&0\\ 0&2b^{2}&0&0\\ 0&0&2b^{2}&-4c\\ 0&0&-4c&4\sin^{4}\theta_{W}\end{pmatrix}$
with $b=(\tan^{2}\theta_{W}\sin^{2}\theta_{W}-\sin\theta_{W}\cos\theta_{W})$ y $c=\sin^{2}\theta_{W}\tan\theta_{W}(\sin^{2}\theta_{W}-\cos^{2}\theta_{W})$
$\Gamma_{AZ}^{J=0}$ $=\dfrac{e^{2}g^{2}}{128\sqrt{3}\pi M_{V}^{2}}\begin{pmatrix}d^{2}&0&0&0\\ 0&d^{2}&0&0\\ 0&0&d^{2}&f\\ 0&0&f&\dfrac{\sin^{2}(2\theta_{W})}{\cos^{2}\theta_{W}}\end{pmatrix}$ (26)
$\Gamma_{AZ}^{J=2}$ $=\dfrac{e^{2}g^{2}}{256\pi M_{V}^{2}}\left(\dfrac{1+3\cos(2\theta)}{2\sqrt{6}}+\sin^{2}\theta\right)\begin{pmatrix}d^{2}&0&0&0\\ 0&d^{2}&0&0\\ 0&0&d^{2}&f\\ 0&0&f&\dfrac{\sin^{2}(2\theta_{W})}{\cos^{2}\theta_{W}}\end{pmatrix}$
with $d=(\cos\theta_{W}-\sin\theta_{W})$ y $f=(\cos\theta_{W}-\sin\theta_{W})^{4}(1+\tan\theta_{W})^{2}$
3.3.2 Higgs Boson-Dark Matter Vector
We follow a similar procedure to take into accounts the Higgs Boson and the Dark Matter interactions. In this case, we require the quartic terms from Lagrangian (2), the quartic terms involving the new vector field and the Higgs boson doublet. The resulting annihilation matrix can be written as follows:
$\Gamma_{Higgs}=\dfrac{1}{32\pi M_{V}^{2}}\begin{pmatrix}\left(\dfrac{3}{2}+\sqrt{2}\right)^{2}\lambda_{4}^{2}&0&0&0\\ 0&\dfrac{9}{4}\lambda_{4}^{2}&0&0\\ 0&0&\dfrac{1}{4}(7\lambda_{2}^{2}+6\lambda_{2}\lambda_{3}+5\lambda_{3}^{2})&\dfrac{1}{2}(\lambda_{3}+2\lambda_{4})^{2}\\ 0&0&\dfrac{1}{2}(\lambda_{3}+2\lambda_{4})^{2}&\dfrac{3}{4}\lambda_{2}^{2}+(\lambda_{2}+\lambda_{3})^{2}\end{pmatrix}$ (27)
3.4 Numerical Results
Now, we have all the needed ingredients to compute the thermal averaged cross-section $\braket{\sigma v}$ using equation (8). When we expand over all the annihilation channels, $\braket{\sigma v}$ takes the form:
$\braket{\sigma v}=\braket{\sigma v}_{W^{+}W^{-}}+\braket{\sigma v}_{ZZ}+\braket{\sigma v}_{\gamma\gamma}+\braket{\sigma v}_{Z\gamma}+\braket{\sigma v}_{hh}.$ (28)
Our results are shown in figures 2 and 3 . The first one shows our predictions without taking into account the non-minimal interactions ($\kappa=0$) while in the second figure we used $\kappa=1$. In both cases, there are several noteworthy characteristics of the Sommerfeld enhancement that merit discussion. Firstly, as usual, the SE is a velocity-dependent phenomenon. For instance, no structure is observed for a relative velocity of $v=0.3$ (dashed lines in figures 2 and 3), which is representative of the early Universe, when the relic density was formed. In contrast, for smaller velocities, like the one expected in the galaxy ($v=10^{-3}$), significant oscillations and resonances appear. We have observed that the small oscillations dependent on interactions mediated by the Higgs boson. Indeed, we see that they are attenuated for large values of $\lambda$ parameter Interestingly, the presence of the non-minimal interactions introduce an additional resonance at a lower mass. However, the $\kappa$-terms don’t modify significantly the size of the oscillations nor the amplitude of the resonances. For masses that exhibit such resonances (roughly around $1.9$ and $6$ TeV), the Sommerfeld factor experiences a significant boost of about 5 order of magnitude, for all the values of portal coupling constant $\lambda$.
Figure 2: Annihilation cross-section (fundamental representation spin-1 of $SU(2)_{L}$) for the process $v_{0}a_{0}\longrightarrow SMSM$ with the Sommerfeld Enhancement factor, considering velocities $v=10^{-3}$ (galactic Halo, shown as continuous lines, green, blue, black and gray lines ) and $v=0.3$ (Relic density, indicated by dashed lines), and different values of the Higgs portal coupling constant $\lambda$, with $\kappa_{1}=\kappa_{2}=0$. The green line indicates that it is the spin-1 triplet representation of $SU(2)_{L}$, while the red line indicates that it is the spin-1/2 doublet representation of $SU(2)_{L}$.
Figure 3: Annihilation cross-section (fundamental representation spin-1 of $SU(2)_{L}$) for the process $v_{0}a_{0}\longrightarrow SMSM$ with the Sommerfeld Enhancement factor, considering velocities $v=10^{-3}$ (galactic Halo, shown as continuous lines, green, blue, black and gray lines ) and $v=0.3$ (Relic density, indicated by dashed lines), and different values of the Higgs portal coupling constant $\lambda$, with $\kappa_{1}=\kappa_{2}=1$. The green line indicates that it is the spin-1 triplet representation of $SU(2)_{L}$, while the red line indicates that it is the spin-1/2 doublet representation of $SU(2)_{L}$.
4 Photons Flux from Annihilation neutral Dark Matter particles
The $\gamma$-ray flux resulting from the annihilation of neutral Dark Matter particles, as described in [17], in a region of the sky with a solid angle $\Omega$ is given by:
$\dfrac{d\Phi}{dE_{\gamma}}=\sum_{i}\dfrac{1}{8\pi}J(\Omega)\dfrac{\braket{\sigma v}_{i}}{M_{V}^{2}}\left(\dfrac{dN_{i}}{dE_{\gamma}}\right)$ (29)
Here, $J(\Omega)$ represents the astrophysical $J$-factor, which measures the relevance and interest of the astrophysical target for Dark Matter annihilation.
[…]
where $\rho[r(\ell,\psi)]$ is the DM density and depends of the assumed astrophysical profile. The $J$-factor for dwarf galaxies is approximately $J\approx 10^{19}$ GeV2/cm5 while for the Andromeda galaxy, $J\approx 10^{20}$ GeV2/cm5, and for the Milky Way’s Galactic Center, $J\approx 10^{25}$ GeV2/cm5. If the solid angle changes by one degree, the $J$-factor undergoes an approximate change of $J\approx 10^{24}$ GeV2/cm5 [21].
The distribution, denoted as $\left(\frac{dN_{i}}{dE_{\gamma}}\right)$, characterizes the number of photons with energy $E_{\gamma}$, while $\braket{\sigma v}_{i}$ represents the averaged annihilation cross-section from the i-th channel of annihilation. This quantity dependents on the branching ratio, and is defined as $\braket{\sigma v}_{\gamma\gamma}=BR_{\gamma\gamma}\times\braket{\sigma v}_{XX^{\prime}}$. Here, $\braket{\sigma v}_{XX^{\prime}}$ corresponds to the total cross-section for Dark Matter annihilation, as shown in equation (8).
Using the expressions above, we calculate the photon flux resulting from the annihilation cross section (8), which serves as a method for the indirect detection of DM. This entails the integration of equation (29). Figures 4 and 5 depict the photon flux from VDM annihilation per second for $\kappa=0$ and $\kappa=1$ respectively. In all the cases we have used $\lambda=0.5$. We present the signals of photon flux resulting from the annihilation of VDM particles from the center of the galaxy for different final states ($\gamma\gamma$, $W^{+}W^{-}$, $ZZ$, and $hh$), with a solid angle of $1^{\circ}$, compared with the canonical cross-section with a mass of 100 GeV. Additionally, we show the total flux predicted at different values of the DM mass. We selected the masses to coincide with peaks and local minimums in the SE factor. We compare our results with Fermi-LAT sesitivity and the projected CTA sensitivity [22].
Figure 4: Photons Flux from Annihilation of neutral Dark Matter particles. Where it is done for the Navarro-Frenk-White (NFW) profile, with a Line of sight (los) angle in longitude (llos) of $1^{\circ}$, and latitude (blos) of $0^{\circ}$. This photon flux is measured at the center of the galaxy, with $\kappa=0$. We show differential sensitivity of CTA [22] and Fermi-Lat [23] in the galactic center.
Figure 5: Photons Flux from Annihilation of neutral Dark Matter particles. Where it is done for the Navarro-Frenk-White (NFW) profile, with a Line of sight (los) angle in longitude (llos) of $1^{\circ}$, and latitude (blos) of $0^{\circ}$. This photon flux is measured at the center of the galaxy, with $\kappa=1$. We show differential sensitivity of CTA [22] and Fermi-Lat [23] in the galactic center.
5 Summary
In this paper, we have studied the Sommerfeld enhancement in a model where the DM candidate is a neutral component of a new massive vector field in the fundamental representation of $SU(2)_{L}$. We found that the annihilation cross-section predicted by our model is generally bigger than the one predicted in the cases of a spin $1/2$ doublet and a spin $1$ triplet. Additionaly, the presence of non-minimal interactions of the doublet with the standard gauge model produces a richer resonance structure. In fact, in the absence of the non-minimal interaction, only one big resonance in $\sigma v$ is observed for $M_{V}\approx 6$ TeV. However, when the non-minimal interactions are present, a second peak appears at $M_{V}\approx 2$ TeV, in the low velocities regime.
As a product of our work, we used our results on $\sigma v$ to predict the photon flux produced by DM annihilation at the center of the galaxy.
Acknowledgments
This work was partially found by Fondecyt grant 1230110, grant PIIC 2021-II UTFSM, ANID BECA/DOCTORADO NACIONAL 21210616, ANID PIA/APOYO AFB220004, and ANID—The Millennium Science Initiative Program ICN2019_044. We are very thankful to Motoko Fujiwara and Patricio Escalona for discussions and suggestions that contributed significantly to improve the quality of this paper.
Appendix A Electroweak Matrix
A.1 Gauge-Dark Matter
In the context of non-relativistic dark matter, we need to determine the potentials that interact with these particles in a low-velocity regime. For this, we need to determine the real potential (cubic terms) of the Lagrangian (2) or effective action of $SU(2)_{L}$, which is the electroweak matrix, where the conserved currents interact with the gauge fields of the Standard Model, resulting in the following expression:
$S_{eff,1}\approx\int d^{4}x~{}\left[J_{\mu}^{W^{+}}(x)W^{+\mu}(x)+J_{\mu}^{W^{-}}(x)W^{-\mu}(x)+J_{\mu}^{Z}(x)Z^{\mu}(x)+J_{\mu}^{A}(x)A^{\mu}(x)\right]$ (31)
After integrating out the gauge bosons in the path integral, we obtain the effective action dependent on the currents. We need to calculate the conserved currents from the cubic terms of the Lagrangian (2); the details of the calculation are in  [4].
$S_{eff,1}=-\int\dfrac{d^{4}xd^{3}y}{8\pi|\vec{x}-\vec{y}|}($ $J_{A}^{0}(x)J_{A}^{0}(x^{0},\vec{y})+J_{Z}^{0}(x)J_{Z}^{0}(x^{0},\vec{y})e^{-M_{Z}|\vec{x}-\vec{y}|}$
$+2J_{W^{+}}^{0}(x)J_{W^{-}}^{0}(x^{0},\vec{y})e^{-M_{W}|\vec{x}-\vec{y}|})$ (32)
It is essential to consider only pairs in the form of vectors, as indicated in (14), while other particle combinations are prohibited. Consequently, the currents in the non-relativistic limit are as follows:
$J_{A}^{0}(x)J_{A}^{0}(x^{0},\vec{y})=-2g^{2}(c_{W}(1-(\kappa_{1}+\kappa{2})-s_{W}t_{W}^{2})^{2}$ $(-[v_{+}^{\dagger}(x)v_{-}^{\dagger}(x^{0},\vec{y})v_{+}(x)v_{-}(x^{0},\vec{y})]$ (33)
$J_{Z}^{0}(x)J_{Z}^{0}(x^{0},\vec{y})$ $=-g^{2}(c_{W}(\kappa_{1}+\kappa_{2})+s_{W})^{2}[v_{+}^{\dagger}(x)v_{-}^{\dagger}(x^{0},\vec{y})v_{+}(x)v_{-}(x^{0},\vec{y})$
$+\left(\dfrac{v_{0}^{\dagger}(x)v_{0}^{\dagger}(x^{0},\vec{y})}{\sqrt{2}}\dfrac{v_{0}(x)v_{0}(x^{0},\vec{y})}{\sqrt{2}}\right)$
$+\left(\dfrac{a_{0}^{\dagger}(x)a_{0}^{\dagger}(x^{0},\vec{y})}{\sqrt{2}}\dfrac{a_{0}(x)a_{0}(x^{0},\vec{y})}{\sqrt{2}}\right)$
$-\left(v_{0}^{\dagger}(x)a_{0}^{\dagger}(x^{0},\vec{y})v_{0}(x)a_{0}(x^{0},\vec{y})\right)]$ (34)
$J_{W_{+}}^{0}(x)J_{W_{-}}^{0}(x^{0},\vec{y})=$ $-\left[\left(\dfrac{3}{8}-2\sqrt{2}\right)(\kappa_{1}+\kappa_{2})+2\sqrt{2}\right]^{2}$
$g^{2}\left[v_{0}^{\dagger}(x)a_{0}^{\dagger}(x^{0},\vec{y})v_{+}(x)v_{-}(x^{0},\vec{y})-v_{-}^{\dagger}(x)v_{+}^{\dagger}(x^{0},\vec{y})a_{0}(x)v_{0}(x^{0},\vec{y})\right]$ (35)
where $s_{W}=\sin\theta_{W}$, $c_{W}=\cos\theta_{W}$ and $t_{W}=\tan\theta_{W}$. Here the terms that are omitted are the ones that do not lead to pairs $v_{0}v_{0}$, $a_{0}a_{0}$, $v_{0}a_{0}$ and $v_{+}v_{-}$. By plugging these currents in equation (A.1), we obtain:
$S_{eff,1_{NR}}=-\int d^{4}xd^{3}y~{}s(x,\vec{y})^{\dagger}V_{g-V}(|\vec{x}-\vec{y}|)s(x,\vec{y})$ (36)
where $V_{g-V}(|\vec{x}-\vec{y}|)$ is the electroweak matrix, with $r=|\vec{x}-\vec{y}|$, therefore:
$V_{g-V}(r)=-\dfrac{g^{2}}{8\pi r}\begin{pmatrix}4a^{2}e^{-M_{Z}r}&0&0&0\\ 0&4a^{2}e^{-M_{Z}r}&0&0\\ 0&0&16a^{2}e^{-M_{Z}r}&b^{2}e^{-M_{W}r}\\ 0&0&b^{2}e^{-M_{W}r}&a^{2}e^{-M_{Z}r}-2(c_{W}(1-(\kappa_{1}+\kappa{2})-s_{W}t_{W}^{2})^{2}\end{pmatrix}$ (37)
with $a=(c_{W}(\kappa_{1}+\kappa_{2})+s_{W})$, $b=(\frac{3}{8}-2\sqrt{2})(\kappa_{1}+\kappa_{2})+2\sqrt{2}$
A.2 Higgs Boson-Dark Matter
We are now employing the Feynman Gauge. The masses of the Goldstone bosons, namely $\phi^{0}$, $\phi^{0\ast}$, and ϕ±superscriptitalic-$\phi^{\pm}$, originating from the doublet (18), are $M_{Z}$ and $M_{W}$. The Higgs particle is significantly lighter than Dark Matter particles. The interacting Lagrangian (2) between the Higgs Boson and the Dark Matter Vector is as follows:
$\mathcal{L}_{h-V}=-\lambda_{2}(\phi^{\dagger}\phi)(V_{\mu}^{\dagger}V^{\mu})-\lambda_{3}(\phi^{\dagger}V_{\mu})(V^{\mu\dagger}\phi)-\dfrac{\lambda_{4}}{2}\left[(\phi^{\dagger}V_{\mu})(\phi^{\dagger}V^{\mu})+(V^{\mu\dagger}\phi)(V_{\mu}^{\dagger}\phi)\right]$ (38)
In order to construct pairs such as $v_{0}v_{0},v_{0}a_{0},a_{0}a_{0}$, and $v_{+}v_{-}$, we need to substitute equations (1) and (18) into (38). Consequently, we obtain:
$V_{h-V}(r)=\dfrac{v^{2}}{32\pi M_{V}^{2}r}\begin{pmatrix}8\lambda_{4}^{2}a&0&0&0\\ 0&8\lambda_{4}^{2}a&0&0\\ 0&0&\dfrac{1}{2}(\lambda_{2}+\lambda_{3})^{2}a&(\lambda_{3}+2\lambda_{4})^{2}e^{-M_{W}r}\\ 0&0&(\lambda_{3}+2\lambda_{4})^{2}e^{-M_{W}r}&\lambda_{2}^{2}a\end{pmatrix}$ (39)
with $a=(2e^{-M_{h}r}+e^{-M_{Z}r})$, this is electroweak matrix (Higgs Boson-Dark Matter interaction)
Appendix B Annihilation Matrix
B.1 Gauge-Dark Matter Interaction
Non-relativistic vector dark matter particles have the potential to annihilate into relativistic particles, specifically gauge bosons of $SU(2)_{L}$. In quantum mechanics (Schrödinger equation), these annihilations are represented by an imaginary potential or an absorptive term (quartic terms) in the action, which explains the disappearance of pairs of non-relativistic particles.
These absorptive terms originate from $S_{eff,2}$. To provide an interpretation, we can visualize $S_{eff,2}$ in terms of Feynman diagrams. In this representation, the matrix $\Delta$ characterizes the propagators of the gauge fields, and each $\mathcal{U}$ represents a vertex featuring two gauge fields and two vector dark matter fields, as described by Chowdhury et al. [4, 24].
Then, the effective action of absorptive terms is:
$S_{eff,2}$ $=-iTr\log\left(\Delta^{-1}\mathcal{U}\Delta^{-1}\mathcal{U}\right)$
$=\frac{i}{2\pi}\int d^{4}x~{}Tr\left[\begin{pmatrix}U_{AA}(x)&U_{AZ}(x)&U_{AW}(x)&U_{AW}^{\ast}(x)\\ U_{AZ}(x)&U_{ZZ}(x)&U_{ZW}(x)&U_{ZW}^{\ast}(x)\\ U_{AW}^{\ast}(x)&U_{ZW}^{\ast}(x)&U_{WW}(x)&0\\ U_{AW}^{\ast}(x)&U_{ZW}^{\ast}(x)&0&U_{WW}^{\ast}(x)\\ \end{pmatrix}\right]^{2}$ (40)
We need only pairs like $V_{0}V_{0},V_{0}^{\ast}V_{0}^{\ast},V_{0}V_{0}^{\ast}$ and $V_{+}V_{-}$, then
$S_{eff,2}$ $=\dfrac{i}{2\pi}\int~{}d^{4}x\left(U_{AA}(x)^{2}+U_{ZZ}(x)^{2}+2U_{WW}(x)^{2}+2U_{AZ}(x)^{2}\right)$
$=-\int~{}d^{4}x\dfrac{g^{2}}{8}[2W_{\mu}^{+}V_{\nu}^{-}V^{+~{}\nu}W^{-~{}\mu}+2W_{\mu}^{-}(V_{\nu}^{0})^{\ast}V^{0~{}\nu}W^{+~{}\mu}$
$+A_{\mu}\left\{\sin^{2}\theta_{W}(1-\tan\theta_{W})(V_{\nu}^{0})^{\ast}V^{0~{}\nu}+2\sin^{2}\theta_{W}V_{\nu}^{-}V^{+~{}\nu}\right\}A^{\mu}$
$+Z_{\mu}\left\{(\tan^{2}\theta_{W}\sin^{2}\theta_{W}-\sin\theta_{W}\cos\theta_{W})(V_{\nu}^{0})^{\ast}V^{0~{}\nu}+2\sin^{2}\theta_{W}V_{\nu}^{-}V^{+~{}\nu}\right\}Z^{\mu}$
$+A_{\mu}((\sin\theta_{W}(2\cos\theta_{W}-\sin\theta_{W}(1+\tan\theta_{W}))(V_{\nu}^{0})^{\ast}V^{0~{}\nu}$
$+\cos(2\theta_{W})\tan\theta_{W}V_{\nu}^{-}V^{+~{}\nu})Z^{\mu}$
$+Z_{\mu}(\sin\theta_{W}(\cos\theta_{W}-\sin\theta_{W})(V_{\nu}^{0})^{\ast}V^{0~{}\nu}$
$+\tan\theta_{W}(\cos^{2}\theta_{W}-\sin^{2}\theta_{W})V_{\nu}^{-}V^{+~{}\nu})A^{\mu}](1+1+1+1)$
$=-\dfrac{g^{2}}{2}[2W_{\mu}^{+}V_{\nu}^{-}V^{+~{}\nu}W^{+~{}\mu}+2W_{\mu}^{-}(V_{\nu}^{0})^{\ast}V^{0~{}\nu}W^{+~{}\mu}$
$+A_{\mu}\left\{\sin^{2}\theta_{W}(1-\tan\theta_{W})(V_{\nu}^{0})^{\ast}V^{0~{}\nu}+2\sin^{2}\theta_{W}V_{\nu}^{-}V^{+~{}\nu}\right\}A^{\mu}$
$+Z_{\mu}\left\{(\tan^{2}\theta_{W}\sin^{2}\theta_{W}-\sin\theta_{W}\cos\theta_{W})(V_{\nu}^{0})^{\ast}V^{0~{}\nu}+2\sin^{2}\theta_{W}V_{\nu}^{-}V^{+~{}\nu}\right\}Z^{\mu}$
$+A_{\mu}((\sin\theta_{W}(2\cos\theta_{W}-\sin\theta_{W}(1+\tan\theta_{W}))(V_{\nu}^{0})^{\ast}V^{0~{}\nu}$
$+\cos(2\theta_{W})\tan\theta_{W}V_{\nu}^{-}V^{+~{}\nu})Z^{\mu}$
$+Z_{\mu}(\sin\theta_{W}(\cos\theta_{W}-\sin\theta_{W})(V_{\nu}^{0})^{\ast}V^{0~{}\nu}$
$+\tan\theta_{W}(\cos^{2}\theta_{W}-\sin^{2}\theta_{W})V_{\nu}^{-}V^{+~{}\nu})A^{\mu}]$ (41)
In terms of s’vectors, the effective action is
$S_{eff,2}=2i\int d^{4}xd^{3}y~{}s(x,\vec{y})^{\dagger}\Gamma_{gauge}\delta^{3}(|x-y|)s(x,\vec{y})$ (42)
The Q’s terms are:
$U_{WW}^{2}(x)$ $=\dfrac{g^{4}}{16}(V_{\nu}^{-}V^{+~{}\nu}+(V_{\nu}^{0})^{\ast}V^{0~{}\nu})^{2}$ (43)
$=\dfrac{g^{4}}{16}(V_{i}^{-}V^{+~{}j}V_{i}^{-}V^{+~{}j}+(V_{i}^{0})^{\ast}V^{0~{}j}(V_{i}^{0})^{\ast}V^{0~{}j}$
$+2(V_{i}^{0})^{\ast}V^{0~{}j}(V_{i}^{-})^{\ast}V^{+~{}j})$
$U_{AA}^{2}(x)=\dfrac{e^{4}}{64}\left[(1-\tan\theta_{W})^{2}((V_{\nu}^{0})^{\ast}V^{0~{}\nu})^{2}+4(V_{\nu}^{-}V^{+~{}\nu})^{2}+4(1-\tan\theta_{W})(V_{\nu}^{0})^{\ast}V^{0~{}\nu}V_{\nu}^{-}V^{+~{}\nu}\right]$ (44)
$U_{ZZ}^{2}(x)=\dfrac{g^{4}}{64}$ $[(\tan^{2}\theta_{W}\sin^{2}\theta_{W}-\sin\theta_{W}\cos\theta_{W})^{2}((V_{\nu}^{0})^{\ast}V^{0~{}\nu})^{2}+4\sin^{4}\theta_{W}(V_{\nu}^{-}V^{+~{}\nu})^{2}$
$-4\sin^{2}\theta_{W}(\tan^{2}\theta_{W}\sin^{2}\theta_{W}-\sin\theta_{W}\cos\theta_{W})^{2}(V_{\nu}^{0})^{\ast}V^{0~{}\nu}V_{\nu}^{-}V^{+~{}\nu}]$ (45)
$U_{AZ}^{2}(x)$ $=\dfrac{e^{2}g^{2}}{64}[(\cos\theta_{W}-\sin\theta_{W})^{2}((V_{\nu}^{0})^{\ast}V^{0~{}\nu})^{2}+\dfrac{\cos^{2}(2\theta_{W})}{\cos^{2}\theta_{W}}{\cos\theta_{W}}(V_{\nu}^{-}V^{+~{}\nu})^{2}$
$+(\cos\theta_{W}-\sin\theta_{W})^{4}(1+\tan\theta_{W})^{2}(V_{\nu}^{0})^{\ast}V^{0~{}\nu}V_{\nu}^{-}V^{+~{}\nu}]$ (46)
However, we need to establish a relationship between polarization vectors and angular momentum, as discussed in [7], where:
$\epsilon_{i}(p_{-})\epsilon_{j}(p_{+})\longrightarrow S_{ij}^{J,J_{z}}$ (47)
The annihilation matrices (Gauge-Dark Matter) are:
$\Gamma_{WW}^{J=0}=\dfrac{2g^{4}}{64\sqrt{3}\pi M_{V}^{2}}\begin{pmatrix}2&0&0&0\\ 0&2&0&0\\ 0&0&1&1\\ 0&0&1&1\end{pmatrix}$ (48)
$\Gamma_{WW}^{J=2}=\dfrac{g^{4}}{64\pi M_{V}^{2}}\left(\dfrac{1+3\cos(2\theta)}{2\sqrt{6}}+\sin^{2}\theta\right)\begin{pmatrix}2&0&0&0\\ 0&2&0&0\\ 0&0&1&1\\ 0&0&1&1\end{pmatrix}$ (49)
$\Gamma_{AA}^{J=0}=\dfrac{e^{4}}{128\sqrt{3}\pi M_{V}^{2}}\begin{pmatrix}2a^{2}&0&0&0\\ 0&2a^{2}&0&0\\ 0&0&a^{2}&4a^{2}\\ 0&0&4a^{2}&4\end{pmatrix}$ (50)
$\Gamma_{AA}^{J=2}=\dfrac{e^{4}}{256\pi M_{V}^{2}}\left(\dfrac{1+3\cos(2\theta)}{2\sqrt{6}}+\sin^{2}\theta\right)\begin{pmatrix}2a^{2}&0&0&0\\ 0&2a^{2}&0&0\\ 0&0&a^{2}&4a^{2}\\ 0&0&4a^{2}&4\end{pmatrix}$ (51)
with $a=(1-\tan\theta_{W})$
$\Gamma_{ZZ}^{J=0}$ $=\dfrac{g^{4}}{128\sqrt{3}\pi M_{V}^{2}}\begin{pmatrix}2b^{2}&0&0&0\\ 0&2b^{2}&0&0\\ 0&0&2b^{2}&-4c\\ 0&0&-4c&4\sin^{4}\theta_{W}\end{pmatrix}$ (52)
$\Gamma_{ZZ}^{J=2}$ $=\dfrac{g^{4}}{256\pi M_{V}^{2}}\left(\dfrac{1+3\cos(2\theta)}{2\sqrt{6}}+\sin^{2}\theta\right)\begin{pmatrix}2b^{2}&0&0&0\\ 0&2b^{2}&0&0\\ 0&0&2b^{2}&-4c\\ 0&0&-4c&4\sin^{4}\theta_{W}\end{pmatrix}$
with $b=(\tan^{2}\theta_{W}\sin^{2}\theta_{W}-\sin\theta_{W}\cos\theta_{W})$ y $c=\sin^{2}\theta_{W}\tan\theta_{W}(\sin^{2}\theta_{W}-\cos^{2}\theta_{W})$
$\Gamma_{AZ}^{J=0}$ $=\dfrac{e^{2}g^{2}}{128\sqrt{3}\pi M_{V}^{2}}\begin{pmatrix}d^{2}&0&0&0\\ 0&d^{2}&0&0\\ 0&0&d^{2}&f\\ 0&0&f&\dfrac{\sin^{2}(2\theta_{W})}{\cos^{2}\theta_{W}}\end{pmatrix}$ (53)
$\Gamma_{AZ}^{J=2}$ $=\dfrac{e^{2}g^{2}}{256\pi M_{V}^{2}}\left(\dfrac{1+3\cos(2\theta)}{2\sqrt{6}}+\sin^{2}\theta\right)\begin{pmatrix}d^{2}&0&0&0\\ 0&d^{2}&0&0\\ 0&0&d^{2}&f\\ 0&0&f&\dfrac{\sin^{2}(2\theta_{W})}{\cos^{2}\theta_{W}}\end{pmatrix}$
with $d=(\cos\theta_{W}-\sin\theta_{W})$ y $f=(\cos\theta_{W}-\sin\theta_{W})^{4}(1+\tan\theta_{W})^{2}$
B.2 Higgs-Dark Matter Interaction
From the Lagrangian (38), we can extract the quartic terms in the effective action. These terms provide the annihilation matrix for the interaction between the Higgs boson and the Dark Matter Vector. We need to operate on the vector in $SU(2)_{L}$ (1) and the Higgs boson doublet (18). Consequently,
$\mathcal{L}_{Quartic~{}h-V}$ $=-\lambda_{2}[\left(\phi^{-}\phi^{+}+\dfrac{h^{2}}{2}-i\dfrac{(\phi^{0\ast}h+h^{\ast}\phi^{0})}{\sqrt{2}}-(\phi^{0})^{2}\right)V_{\mu}^{-}V^{+\mu}$ (54)
$+\left(\phi^{-}\phi^{+}+\dfrac{h^{2}}{2}-i\dfrac{(\phi^{0\ast}h+h^{\ast}\phi^{0})}{\sqrt{2}}-(\phi^{0})^{2}\right)V_{\mu}^{0\ast}V^{0\mu}]$
$-\lambda_{3}[\phi^{-}\phi^{+}V_{\mu}^{+}V^{-\mu}+\dfrac{1}{\sqrt{2}}(i\phi^{0}+h)\phi^{-}V_{\mu}^{+}V^{0\mu\ast}+\dfrac{1}{\sqrt{2}}(-i\phi^{0\ast}+h^{\ast})\phi^{+}V_{\mu}^{0}V^{-\mu}$
$+\left(\dfrac{h^{2}}{\sqrt{2}}-(\phi^{0})^{2}\right)V_{\mu}^{0}V^{0\mu\ast}]$
$+\left(\dfrac{h^{2}}{\sqrt{2}}-(\phi^{0})^{2}\right)V_{\mu}^{0}V^{0\mu\ast}]$
$-\lambda_{4}[\left(-i\phi^{0\ast}+\dfrac{h^{\ast}}{\sqrt{2}}\right)\phi^{-}V^{\mu+}V^{0\mu}+\left(-i\phi^{0\ast}+\dfrac{h^{\ast}}{\sqrt{2}}\right)\phi^{-}V^{\mu 0}V^{+\mu}$
$+\left(\dfrac{h^{\ast}}{2}-i\sqrt{2}h^{\ast}\phi^{0\ast}+(\phi^{0\ast})^{2}\right)V_{\mu}^{0}V^{0\mu}+\left(i\phi^{0}+\dfrac{h}{\sqrt{2}}\right)\phi^{+}V^{-\mu}V_{\mu}^{0\ast}$
$+\left(i\phi^{0}+\dfrac{h}{\sqrt{2}}\right)\phi^{+}V^{0\mu\ast}V_{\mu}^{-}+\left(-(\phi^{0})^{2}+\dfrac{h^{2}}{2}\right)V^{0\mu\ast}V_{\mu}^{0\ast}]$
The permissible terms are $V_{0}V_{0},V_{0}V_{0}^{\ast},V_{0}^{\ast}V_{0}^{\ast}$, and $V_{+}V_{-}$, as we require neutral terms. The resulting annihilation matrix is as follows:
$\Gamma_{h-V}=\dfrac{1}{32\pi M_{V}^{2}}\begin{pmatrix}\left(\dfrac{3}{2}+\sqrt{2}\right)^{2}\lambda_{4}^{2}&0&0&0\\ 0&\dfrac{9}{4}\lambda_{4}^{2}&0&0\\ 0&0&\dfrac{1}{4}(7\lambda_{2}^{2}+6\lambda_{2}\lambda_{3}+5\lambda_{3}^{2})&\dfrac{1}{2}(\lambda_{3}+2\lambda_{4})^{2}\\ 0&0&\dfrac{1}{2}(\lambda_{3}+2\lambda_{4})^{2}&\dfrac{3}{4}\lambda_{2}^{2}+(\lambda_{2}+\lambda_{3})^{2}\end{pmatrix}$ (55)